The December 31, 2019, balance sheet for Zachary Corporation is presented here. These are the only accounts on Zachary’s balance sheet. Amounts indicated by question marks (?) can be calculated using the following additional information:

ZACHARY CORPORATION
Balance Sheet As of December 31, 2019
Assets
Cash $ 32,000
Accounts receivable (net) ?
Inventory ?
Property, plant, and equipment (net) 300,000
$ 450,000
Liabilities and Stockholders’ Equity
Accounts payable (trade) $ ?
Income taxes payable (current) 32,000
Long-term debt ?
Common stock 306,000
Retained earnings ?
$ ?
Additional Information
Current ratio (at year end) 1.5 to 1.0
Total liabilities ÷ Total stockholders’ equity 80 %
Gross margin percentage 30 %
Inventory turnover (Cost of goods sold ÷ Ending inventory) 11.2 times
Gross margin for 2019 $ 375,000

Required

  1. Compute the balance in trade accounts payable as of December 31, 2019.
  2. Compute the balance in retained earnings as of December 31, 2019.
  3. Compute the balance in the inventory account as of December 31, 2019. (Assume that the level of inventory did not change from last year.)

(For all requirements, negative amounts should be indicated by a minus sign. Do not round intermediate calculations. Round your answers to the nearest whole dollar amount.)

a. Accounts payable
b. Retained earnings
c. Inventory
Answer #1

Sales = $375,000/30% = $1,250,000

Cost of goods sold = $1,250,000 - 375,000 = $875,000

Inventory turnover = 11.2 times

Cost of goods sold / Ending inventory = 11.2

$875,000 / Ending inventory = 11.2

Ending inventory = $875,000/11.2 = $78,125

Accounts receivable = $450,000-300,000-78,125-32,000 = $39,875

Current ratio = Current assets / Current liabilities

1.5 = $150,000 / Current liabilities

Current liabilities = $150,000/1.5 = $100,000

Current liabilities = Accounts payable+Income tax payable

Accounts payable = $100,000-32,000 = $68,000

Let total liabilities be x:

Total liabilities = 0.8 stockholders equity

Total stockholders equity = Total assets - Total liabilities

x = 0.8 ( $450,000 - x)

x = $360,000 - 0.8x

x+0.8x = $360,000

x = $360,000/1.8 = $200,000

Total liabilities = $200,000

Total stockholders equity = $450,000 - 200,000 = $250,000

Total stockholders equity = Common stock + Retained earnings

Retained earnings = Total stockholders equity - Common stock

Retained earnings = $250,000 - 306,000 = -$56,000